Soybean meal is 12% protien; cornmeal is 6% protein. How many pounds of each should be mixed together in order to get 240 lb mixture that is 10% protein
How many pounds of the cornmeal should be in the mixture?____lb
How many pounds of the soybean meal should be in the mixture?____lb

Answer by stanbon(57226)

Soybean meal is 12% protien; cornmeal is 6% protein. How many pounds of each should be mixed together in order to get 240 lb mixture that is 10% protein
How many pounds of the cornmeal should be in the mixture?____lb
How many pounds of the soybean meal should be in the mixture?____lb
-------------------
Equation:
protein + protein = protein
0.12x + 0.06*(240-x) = 0.10*240
Multiply thru by 100 to get:
12x + 6*240 - 6x = 10*240
6x = 4*240
x = 160 lb (amount of cornmeal needed in the mixture)
240-x = 80 lb (amount of soybean meal needed in the mixture)
================================================================
Cheers,
Stan H.